# Let LLMs Break Free from Overthinking via Self-Braking Tuning

### v2

@@ 1 Introduction
Building upon the high-quality OpenR1-Math [1] dataset of reasoning trajectories, we constructed two specialized training datasets using these strategies: OpenR1-Math-SBT-E and OpenR1-Math-SBT-D. To further enhance the model’s self-awareness of its reasoning state, we introduce braking prompts at the identified braking points, explicitly simulating the recognition of having sufficiently completed the reasoning process. These prompts enable the model to naturally express an awareness of having reached adequate reasoning depth, thus promoting autonomous termination without the need for external signals. Our experimental results demonstrate consistent improvements in reasoning efficiency across mathematical benchmarks of varying difficulty. While maintaining high accuracy, token consumption was reduced by 30% to 60%. Our contributions can be summarized as follows:
@@ 2 Methods
Figure 2: Overview of Self-Braking Tuning. (a) Data construction process with overthinking identification and self braking truncation strategies. (b) An example of automatic reasoning termination in a trained Self-Braking LLM.
@@ 2.2 Overthinking Identification
$\kappa_{t}=\frac{1}{TT}\sum_{i=1}^{TT}\mathbb{I}[w_{i}\in\mathcal{M}],\quad\mathbb{I}[w_{i}\in\mathcal{M}]=\begin{cases}1,&\text{if }w_{i}\in\mathcal{M}\\ 0,&\text{otherwise}\end{cases}\$ (2)
where $\mathcal{M}$ represents our curated lexicon of overthinking marker terms (complete list in Appendix B), TT (Total Tokens) represents the total number of tokens in the thinking segment, $\mathbb{I}[\cdot]$ is the indicator function that equals 1 when $w_{i}$ belongs to $\mathcal{M}$ and 0 otherwise.
@@
Overthink $\text{Overthink\hskip 3.30002ptScore}=\beta\times\kappa_{t}+(1-\beta)\times(1-\eta_{s})$ (3)
@@ 4.2 Preserved Reasoning and Redundancy Masking Trade-off
4.4 Natural Language Guidance vs. Special Token Guidance
A fundamental aspect of Self-Braking Tuning is the mechanism used to signal reasoning termination. We compared our natural language guidance approach (epiphany sentences like “I’ve verified my answer, no need to continue…”) with a special token approach using <stop_overthinking> as an explicit control signal. As shown in Table 5, natural language guidance demonstrates superior performance, achieving equivalent accuracy with significantly fewer tokens (1682 vs. 1797). This suggests that metacognitive self-reflection embedded in natural language provides more effective learning signals than explicit control tokens.
This insight aligns with the inherent nature of LRMs, which are fundamentally language models trained to understand semantic and contextual relationships. Natural language guidance leverages the model’s existing capabilities to recognize logical transitions and inference completion, rather than introducing artificial control mechanisms that require learning new conventions. This finding suggests that future approaches to controlling LRM behavior may benefit from working with, rather than around, their foundational linguistic capabilities.
5 Related Works
Large Reasoning Models
Large reasoning models (LRMs) extend traditional language models with advanced reasoning capabilities, often enabled by reinforcement learning. OpenAI’s o1 series [3] marked a key milestone, followed by DeepSeek-R1 [4], which matched o1’s performance through a cost-efficient combination of supervised fine-tuning and RL. Subsequent models like Kimi-k1.5 [7] and QwQ-32B [5] further solidified the LRM era. Concurrently, alternative approaches have emerged to reduce reliance on RL, leveraging supervised fine-tuning and data distillation. Models such as DeepSeek Distill, OpenR1-Math-7B [28], Sky-T1 [29], and LIMO [30] have shown that strong reasoning performance can also be achieved without RL, broadening the design space for LRMs.
Efficient Reasoning
@@ Appendix A Limitations
Appendix B Overthink Markers
@@
Appendix C Algorithms of Self-Braking Tuning
@@
[…]
@@
0:  Reasoning trajectory $T$ with steps $[S_{1},S_{2},...,S_{n}]$, thresholds $\tau_{1}$ and $\tau_{2}$
@@
[…]
@@
[…]
@@
[…]
@@
Appendix D Detailed Experimental Results
@@
Table 6: Performance with varying overthink score thresholds. A lower threshold (e.g., 0.2), which identifies more reasoning instances as overthinking, yields the best overall performance—achieving up to 49% token reduction while preserving 97.4% of baseline accuracy, especially for SBT-Exact. This supports the effectiveness of aggressive pruning in eliminating redundant reasoning across tasks. On easier datasets like GSM8K, accuracy remains stable across all thresholds. For more challenging datasets such as AIME, although 0.2 still performs best overall, slightly higher thresholds (0.3–0.4) show marginal improvements in accuracy at the cost of increased token usage, suggesting that task complexity can influence the sensitivity to pruning aggressiveness.
@@
Guiding Mode GSM8K MATH500 AIME AMC AVERAGE Acc #Tok Acc #Tok Acc #Tok Acc #Tok Acc #Tok Baseline 85.00 514 80.25 1712 16.25 7381 55.94 3503 59.36 3277 Natural Language 84.87 414 77.30 1046 14.17 3381 50.31 1888 56.66 1682 Special Token 84.94 413 77.92 1120 12.34 3647 51.25 2007 56.61 1797
Table 9: Performance comparison between natural language and special token guiding modes for reasoning termination. Natural language guidance achieves comparable or better accuracy with consistently lower token usage across datasets—e.g., 3381 vs. 3647 tokens on AIME, and 1888 vs. 2007 on AMC—demonstrating more efficient reasoning control. Accuracy differences are minor on GSM8K and MATH500, but natural language guidance significantly outperforms special tokens on AIME (14.17% vs. 12.34%). These results suggest that semantically aligned, self-reflective cues are more effective than explicit control tokens, especially in complex reasoning scenarios.

### v4

# Let LRMs Break Free from Overthinking via Self-Braking Tuning

@@ 1 Introduction
Building upon the high-quality OpenR1-Math [1] dataset of reasoning trajectories, we constructed two specialized training datasets using these strategies: OpenR1-Math-SBT-E and OpenR1-Math-SBT-D. To further enhance the model’s self-awareness of its reasoning state, we introduce braking prompts at the identified braking points, explicitly simulating the recognition of having sufficiently completed the reasoning process. These prompts enable the model to naturally express an awareness of having reached adequate reasoning depth, thus promoting autonomous termination without the need for external signals. Our experimental results demonstrate consistent improvements in reasoning efficiency across mathematical benchmarks of varying difficulty. While maintaining high accuracy, token consumption was reduced by 30% to 60%.
Our contributions can be summarized as follows:
@@
2 Related Works
Large reasoning models
Large reasoning models (LRMs) extend traditional language models with advanced reasoning capabilities, often enabled by reinforcement learning. OpenAI’s o1 series [3] marked a key milestone, followed by DeepSeek-R1 [4], which matched o1’s performance through a cost-efficient combination of supervised fine-tuning and RL. Subsequent models like Kimi-k1.5 [7] and QwQ-32B [5] further solidified the LRM era. Concurrently, alternative approaches have emerged to reduce reliance on RL, leveraging supervised fine-tuning and data distillation. Models such as DeepSeek-Distill [4], OpenR1-Math-7B [23], Sky-T1 [24], and LIMO [25] have shown that strong reasoning performance can also be achieved without RL, broadening the design space for LRMs. In addition, recent algorithmic advances specifically target robustness and scalability of reasoning [26, 27, 28]: S3c-Math [29] introduces spontaneous step-level self-correction to let models detect and fix erroneous intermediate steps during chain-of-thought generation, while InftyThink [30] breaks long-context limits by turning monolithic proofs into iterative short-segment reasoning with concise progress summaries, enabling effectively unbounded reasoning depth with bounded computation.
Efficient reasoning
@@ 3 Methods
Figure 2: Overview of Self-Braking Tuning. Left: Data construction process with overthinking identification and self braking truncation strategies. Right: An example of automatic reasoning termination in a trained Self-Braking LLM.
@@ 3.2 Overthinking identification
$\kappa_{t}=\frac{1}{TT}\sum_{i=1}^{TT}\mathbb{I}[w_{i}\in\mathcal{M}],\quad\mathbb{I}[w_{i}\in\mathcal{M}]=\begin{cases}1,&\text{if }w_{i}\in\mathcal{M}\\ 0,&\text{otherwise}\end{cases}\\$ (2)
where $\mathcal{M}$ represents our curated lexicon of overthinking marker terms (complete list in Appendix E), TT (Total Tokens) represents the total number of tokens in the thinking segment, $\mathbb{I}[\cdot]$ is the indicator function that equals 1 when $w_{i}$ belongs to $\mathcal{M}$ and 0 otherwise.
@@
$\text{Overthink\hskip 3.30002ptScore}=\beta\times\kappa_{t}+(1-\beta)\times(1-\eta_{s})$ (3)
@@ 5.2 Preserved reasoning and redundancy masking trade-off
5.3 With vs. without masked redundant thinking
Table 4: MRT comparison. Detailed results shown in Appendix Table 14.
Config. Acc #Tok
Baseline 59.36 3277
w/ MRT 57.83 1673
w/o MRT 58.02 2306
To assess masked redundant thinking (MRT), we compare SBT-E with and without masked segments. In the ablation variant, trajectories truncate directly without retaining masked content.
Results show removing MRT yields minimal accuracy gain (+0.19%) but substantial efficiency loss (+37.8% tokens). Without exposure to masked overthinking patterns, models fail to internalize what constitutes redundancy. Direct truncation provides only implicit stopping signals—models see where reasoning ends but not why. Masked segments serve as explicit negative examples: models observe overthinking without gradient reinforcement, enabling discriminative learning of termination boundaries. This exposure-without-reinforcement principle proves essential for self-regulation while preserving reasoning quality.
5.4 Impact of $\beta$ on RER-OMR balance
Table 5: Performance with different $\beta$ values for overthink score weighting. Detailed results shown in Appendix Table 15.
Method $\boldsymbol{\beta}$ Acc #Tok
Baseline 59.36 3277
SBT-E 0.05 56.48 1762
0.1 57.83 1673
0.15 56.52 1874
0.2 55.86 1809
SBT-D 0.05 56.24 1678
0.1 56.66 1682
0.15 56.21 1784
0.2 55.74 1814
The overthink score (Equation 3) combines structural efficiency $(1-\eta_{s})$ and linguistic markers $\kappa_{t}$ through weighting parameter $\beta$. Table 5 examines sensitivity across $\beta\in\{0.05,0.1,0.15,0.2\}$.
Results show $\beta=0.1$ consistently achieves optimal performance for both SBT variants, with 57.83% accuracy (SBT-E) and 56.66% (SBT-D) while maintaining strong efficiency (1673 and 1682 tokens). Lower values ($\beta=0.05$) over-emphasize linguistic markers, degrading accuracy by 0.4–1.4 points despite minimal token savings. Higher values ($\beta\geq 0.15$) reduce linguistic contribution, causing accuracy drops of 1.3–2.0 points with increased token consumption. This validates prioritizing structural efficiency (90%) while retaining linguistic signals (10%) for robust detection without prompt sensitivity.
@@ 5.4 Impact of $\beta$ on RER-OMR balance
No Guidance 56.39 1801
5.6 Natural language guidance vs. alternative approaches
A fundamental aspect of Self-Braking Tuning is the mechanism used to signal reasoning termination. We compared our natural language guidance approach (epiphany sentences like “I’ve verified my answer, no need to continue…”) with both a special token approach using <stop_overthinking> and a configuration without any explicit guidance.
Table 7 shows natural language guidance achieves optimal performance (56.66% accuracy, 1682 tokens). The necessity of explicit termination signals is confirmed by the no-guidance ablation, which degrades both accuracy (0.27-0.27%) and efficiency (+7.1% tokens). Compared to special tokens (56.61%, 1797 tokens), natural language guidance maintains comparable accuracy with 6.4% fewer tokens by leveraging models’ existing semantic understanding of logical transitions rather than requiring learning of artificial control conventions.
@@ 6 Conclusion
Acknowledgement
This work is supported by the National Natural Science Foundation of China (No. 62376245), National Key Research and Development Project (No. 2024YFB3312900), the Key Research and Development Program of Zhejiang Province, China (No. 2024C03255), the Fundamental Research Funds for the Central Universities (226-2024-00170), MOE Engineering Research Center of Digital Library, CCF-Tencent Rhino-Bird Open Research Fund, and ZJU Kunpeng&Ascend Center of Excellence.
@@ Appendix A Limitations
Appendix B Autonomous reasoning regulation in self-braking models
A critical distinction of our Self-Braking Tuning framework is that SBT-E and SBT-D are data construction strategies, not inference-time control mechanisms. After training, models exhibit autonomous reasoning regulation without external intervention, dynamically adapting reasoning depth to problem complexity.
B.1 Emergent self-regulation behavior
To validate that self-braking emerges naturally during inference, we analyzed reasoning patterns of Qwen2.5-Math-7B-Instruct-SBT-E on AIME. Models spontaneously generate epiphany sentences (e.g., “Wait, I’ve verified my answer. No need to continue thinking.”) and terminate without external signals.
Table 8: Analysis of autonomous early termination in SBT-trained models on AIME. Models naturally terminate reasoning in about 50% of cases, achieving both higher accuracy and 48–51% token reduction.
Dataset Exit Type % of Cases Acc #Tok
AIME24 Early Exit 50.83% 41.80% 5,692
No Early Exit 49.17% 38.98% 11,084
AIME25 Early Exit 49.17% 41.53% 6,483
No Early Exit 50.83% 32.79% 12,201
Table 8 shows that self-braking is not enforced through hard constraints. Cases with early termination achieve higher accuracy while using fewer tokens—indicating models learn to recognize when additional reasoning becomes counterproductive.
B.2 Problem-adaptive reasoning depth
SBT-trained models autonomously adjust reasoning length based on task difficulty:
[…]
Dataset Difficulty Avg Steps
GSM8K Easy 27.78
MATH500 Medium 51.32
AMC23 Hard 106.40
AIME25 Very Hard 202.23
The substantial variation (27.78 to 202.23 steps) occurs without external difficulty indicators or task-specific prompting, suggesting internalized problem complexity assessment.
B.3 Interpretation of self-regulation mechanisms
The training process exposes models to three key signals: (1) preserved reasoning demonstrating sufficient problem-solving depth, (2) natural language cues marking reasoning completion, and (3) masked redundant segments that models observe but do not learn to reproduce. This combination appears to develop an internal representation of “reasoning sufficiency” that activates during generation.
We hypothesize that the natural language braking prompts (e.g., “Wait, I’ve gotten the same answer multiple times”) serve as anchor points in the representation space, allowing models to recognize similar reasoning states during inference and trigger corresponding termination behavior. The consistency of early termination rates ($\sim$50% on AIME) across individual problems suggests this behavior generalizes beyond memorization of training examples.
Appendix C Out-of-domain generalization
To evaluate whether Self-Braking Tuning generalizes beyond mathematical reasoning, we test SBT-trained models on two out-of-domain benchmarks covering diverse knowledge domains: MMLU-Redux (general knowledge and reasoning) [49] and GPQA-Diamond (graduate-level science questions) [50]. All models were trained exclusively on math data (OpenR1-Math), with no domain-specific fine-tuning for these benchmarks.
Table 10: Out-of-domain evaluation on MMLU-Redux and GPQA-Diamond. Models trained exclusively on mathematical reasoning data demonstrate consistent efficiency gains (26–65% token reduction) across non-mathematical tasks, with accuracy drops typically limited to 1–3%.
Base Model Method MMLU-Redux GPQA-Diamond Average
Acc #Tok Acc #Tok Acc #Tok
Qwen2.5-Math-1.5B Baseline 45.84 2,061 24.75 5,485 35.30 3,773
SBT-E 43.12 1,403 25.06 3,194 34.09 2,299
SBT-D 43.28 1,566 26.20 3,002 34.74 2,284
Qwen2.5-Math-7B Baseline 67.04 3,229 41.15 8,892 54.10 6,061
SBT-E 65.84 1,927 40.40 6,205 53.12 4,066
SBT-D 66.39 1,998 41.29 6,706 53.84 4,352
Llama-3.2-1B Baseline 35.62 1,933 17.99 9,321 26.81 5,627
SBT-E 32.24 770 24.24 3,516 28.24 2,143
SBT-D 33.12 725 23.48 3,157 28.30 1,941
Llama-3.1-8B Baseline 80.53 2,481 37.31 8,918 58.92 5,699
SBT-E 77.46 1,646 36.30 5,346 56.88 3,496
SBT-D 77.74 1,668 36.91 6,717 57.33 4,192
Cross-domain transferability. SBT achieves 26–65% token reduction across non-mathematical benchmarks despite training exclusively on math data, with minimal accuracy degradation (typically 1–3%). This demonstrates that self-regulation learned from mathematical reasoning generalizes to diverse knowledge domains.
Model size effects. Efficiency gains scale inversely with model size: smaller models (Llama-3.2-1B) achieve 65% token reduction, while larger models (Llama-3.1-8B, Qwen2.5-Math-7B) show 26–40% reductions. Notably, Llama-3.2-1B with SBT-E improves GPQA-Diamond accuracy by 6.25 percentage points (17.99% → 24.24%), indicating that overthinking mitigation can enhance performance on challenging out-of-domain tasks for smaller models.
Architecture consistency. Both domain-specialized (Qwen2.5-Math) and general-purpose (Llama) models benefit comparably from SBT on out-of-domain evaluation, confirming that self-braking represents a transferable meta-cognitive capability rather than task-specific adaptation.
These results establish that reasoning efficiency patterns learned from mathematics transfer broadly to general knowledge and scientific reasoning, supporting SBT’s applicability beyond its training domain.
Appendix D Extended threshold analysis
To comprehensively evaluate the robustness of our overthinking detection mechanism, we conducted an extended threshold analysis spanning $\tau_{1}\in\{0.05,0.1,0.2,0.3,0.4,0.5\}$. This analysis examines how threshold selection affects both the proportion of samples classified as overthinking and the resulting model performance.
Table 11: Extended threshold analysis for overthinking score across $\tau_{1}\in[0.05,0.5]$. Results Averaged across GSM8K, MATH500, AIME, and AMC23. Overthink % indicates the proportion of samples classified as containing overthinking at each threshold.
Method Threshold Overthink % Accuracy #Tokens Reduction
Baseline 59.36 3,277
SBT-E 0.05 75.20 55.14 1,407 57.1%
0.1 65.40 55.59 1,427 56.5%
0.2 60.30 57.83 1,673 48.9%
0.3 50.20 56.70 1,755 46.4%
0.4 41.00 57.38 1,834 44.0%
0.5 2.06 57.12 2,602 20.6%
SBT-D 0.05 74.20 54.90 1,215 62.9%
0.1 62.30 55.90 1,251 61.8%
0.2 62.50 56.66 1,682 48.7%
0.3 50.90 57.47 1,917 41.5%
0.4 40.10 57.36 1,902 42.0%
0.5 0.19 57.09 2,696 17.7%
D.1 Threshold regime characterization
The extended analysis reveals three distinct performance regimes:
Aggressive pruning ($\tau_{1}<0.2$). Low thresholds classify 62–75% of samples as overthinking, achieving maximum token reduction (57–63%) but incurring significant accuracy degradation (2.5–4.5 percentage points). At $\tau_{1}=0.05$, the framework removes reasoning content too aggressively, truncating even necessary exploration steps in complex problems. This regime prioritizes efficiency over reasoning completeness.
Balanced operation ($\tau_{1}\in[0.2,0.4]$). This range identifies 40–62% of samples as overthinking and demonstrates stable performance with <<1.2% accuracy variation while maintaining 41–49% token reduction. The relative insensitivity to exact threshold values within this range indicates robust overthinking detection. Peak performance occurs at $\tau_{1}=0.2$ for both SBT variants, achieving optimal accuracy-efficiency trade-offs (57.83% / 48.9% reduction for SBT-E; 56.66% / 48.7% for SBT-D).
Conservative pruning ($\tau_{1}=0.5$). High thresholds classify only 0.2–2% of samples as overthinking, causing the framework to degenerate toward baseline behavior. Token reduction drops to 18–21% with no compensatory accuracy gains, indicating that insufficient overthinking mitigation fails to realize efficiency benefits.
D.2 Threshold selection rationale
We select $\tau_{1}=0.2$ as the default threshold based on three considerations:
Performance optimality. This threshold achieves the best accuracy within the balanced regime while maintaining substantial efficiency gains across both SBT variants and all evaluated benchmarks.
Empirical stability. The plateau in the $[0.2,0.4]$ range demonstrates that performance degrades gradually rather than catastrophically under threshold perturbations, supporting practical deployability.
Dataset composition. A threshold of 0.2 creates approximately 60% SBT-processed and 40% original trajectories, enabling models to learn both when to terminate (from processed examples) and when to continue reasoning (from preserved examples). This balanced exposure prevents over-generalization of braking behavior while establishing clear termination patterns.
These findings establish that SBT’s effectiveness stems from systematic overthinking identification rather than aggressive truncation, with the optimal operating point balancing reasoning preservation against redundancy elimination.
Appendix E Overthink markers
@@
Appendix F Algorithms of self-braking tuning
@@
7:  $T^{\prime}\leftarrow\text{PreservedSegment}+\text{MaskedSegment}$ {With loss masking on MaskedSegment}
@@
0:  Reasoning trajectory $T$ with steps $S_{1},S_{2},...,S_{n}]$, thresholds $\tau_{1}$ and $\tau_{2}$
@@
1:  $\text{PreservedThinking}\leftarrow\text{Foundation Solution from }T$
2:  $i\leftarrow\text{index of first step after Foundation Solution}$
3:  while $i\leq n$ $\text{CalculateOverthinkScore}(\text{PreservedThinking}+S_{i})<\tau_{1}$ do
4:   $\text{PreservedThinking}\leftarrow\text{PreservedThinking}+S_{i}$
@@
8:  while $i\leq n$ $\text{CalculateOverthinkScore}(\text{PreservedThinking}+\text{MaskedThinking}+S_{i})<\tau_{2}$ do
@@
12:  $T^{\prime}\leftarrow\text{PreservedThinking}+\text{MaskedThinking}$ {With loss masking on MaskedThinking}
@@
Appendix G Detailed experimental results
@@
Table 12: Performance with varying overthink score thresholds. Lower thresholds (0.2) achieve optimal efficiency-accuracy trade-offs—up to 49% token reduction while preserving 97.4% baseline accuracy—validating aggressive pruning effectiveness. Accuracy remains stable across thresholds on simple tasks (GSM8K), while complex tasks (AIME) show marginal accuracy gains at higher thresholds (0.3–0.4) with increased token costs, indicating task-dependent sensitivity to pruning aggressiveness.
@@
Table 14: Performance comparison of Masked Redundant Thinking mechanism across benchmarks of varying difficulty. Removing MRT yields marginal Average accuracy gain (+0.19%) but substantial token increase (+37.8%), with dataset-dependent effects: simple tasks (GSM8K: +9.6% tokens) show minimal impact, while complex reasoning tasks experience severe efficiency degradation (AIME: +46.5% tokens, AMC: +36.4% tokens). These results confirm MRT enables efficient termination learning through exposure-without-reinforcement, with benefits scaling with task complexity.
Configuration GSM8K MATH500 AIME AMC Average Acc #Tok Acc #Tok Acc #Tok Acc #Tok Acc #Tok Baseline 85.00 514 80.25 1712 16.25 7381 55.94 3503 59.36 3277 w/ MRT 84.85 426 77.10 1121 13.75 3101 55.63 2044 57.83 1673 w/o MRT 85.05 467 78.30 1425 14.38 4544 54.37 2788 58.02 2306
Table 15: Performance with different $\beta$ values for overthink score weighting across benchmarks. While $\beta=0.1$ achieves best Average performance for both variants (57.83% SBT-E, 56.66% SBT-D), optimal values vary by dataset: $\beta=0.05$ maximizes accuracy on simple tasks (GSM8K) and minimizes tokens on MATH500, while $\beta=0.1$ dominates on complex tasks (AIME, AMC23). Lower $\beta$ over-emphasizes linguistic markers with limited efficiency gains, while higher $\beta$ degrades both accuracy and efficiency. This validates $\beta=0.1$ as the robust default, balancing structural efficiency (90%) and linguistic signals (10%) across task complexities.
Method $\boldsymbol{\beta}$ GSM8K MATH500 AIME AMC23 Average Acc #Tok Acc #Tok Acc #Tok Acc #Tok Acc #Tok Baseline 85.00 514 80.25 1712 16.25 7381 55.94 3503 59.36 3277 SBT-E 0.05 84.92 433 76.45 1089 13.13 3198 51.42 2328 56.48 1762 0.1 84.85 426 77.10 1121 13.75 3101 55.63 2044 57.83 1673 0.15 84.78 448 76.80 1243 13.54 3487 50.96 2318 56.52 1874 0.2 84.61 441 76.05 1198 12.92 3421 49.86 2176 55.86 1809 SBT-D 0.05 84.89 421 76.50 1087 13.33 3289 50.24 1915 56.24 1678 0.1 84.87 414 77.30 1046 14.17 3381 50.31 1888 56.66 1682 0.15 84.83 437 76.95 1178 13.75 3598 49.31 1923 56.21 1784 0.2 84.71 429 76.30 1154 13.13 3701 48.82 1972 55.74 1814
@@
Table 17: Performance comparison of different guiding mechanisms for reasoning termination. Natural language guidance consistently achieves the best efficiency-accuracy trade-off across benchmarks, with 3381 vs. 3647 tokens on AIME and 1888 vs. 2007 on AMC compared to special tokens. The ablation without any guidance mechanism (No Guidance) demonstrates the necessity of explicit termination signals, showing degraded performance with both reduced accuracy (0.27-0.27%) and increased token consumption (+7.1+7.1%) compared to natural language guidance. These results confirm that semantically aligned, self-reflective cues enable more effective reasoning regulation than explicit control tokens or implicit learning alone.
Guiding Mode GSM8K MATH500 AIME AMC Average Acc #Tok Acc #Tok Acc #Tok Acc #Tok Acc #Tok Baseline 85.00 514 80.25 1712 16.25 7381 55.94 3503 59.36 3277 Natural Language 84.87 414 77.30 1046 14.17 3381 50.31 1888 56.66 1682 Special Token 84.94 413 77.92 1120 12.34 3647 51.25 2007 56.61 1797 No Guidance 84.73 421 77.05 1098 13.54 3523 50.24 2162 56.39 1801